### v2

# Proposal for an experiment to verify Wigner’s rotation at non-relativistic speeds with massive spin-$1/2$ particles

Abstract
The Wigner rotation of quantum particles with spin is one of the fascinating consequences of interplay between two fundamental aspects of Nature: special relativity and quantum mechanics. A direct high precision verification of the Wigner rotation angle of spin-$1/2$ particles would test on the one hand the Lorentz symmetry of Nature and on the other its realization in quantum mechanics. In this paper we propose such an experiment to directly verify Wigner’s rotation in the regime of non-relativistic velocities at $2\cdot 10^{3}$ m/s for massive spin-$1/2$ particles. We discuss how the experiment could be carried out in a laboratory using slow neutrons. The measurement at non-relativistic velocities becomes possible through letting neutrons propagate for a sufficiently long time because Wigner rotation is a cumulative effect.
I Introduction
Over the years, low-energy neutrons have proven to be invaluable in a range of experiments that probe foundational issues in physics Rauch and Werner (2000); Sponar et al. (2010a); Hasegawa et al. (2006); Bartosik et al. (2009); Klepp et al. (2014); Sponar et al. (2021). They are useful test particles in interferometric, spectroscopic and scattering experiments probing low-energy scales and are a complementary tool to particle colliders. Their significance lies in providing direct tests of fundamental concepts to a high degree of accuracy. One area they might be usefully deployed is in precision measurements of the Wigner rotation, which first came to the fore in the context of atomic physics when Thomas showed that the hydrogen atom needs a relativistic correction due to a phenomenon that came to be called the Thomas precession Thomas (1926). In a nutshell, the Wigner rotation is a phenomenon that if a physical system undergoes non-collinear Lorentz boosts, then it is rotated relative to the original frame. Wigner’s rotation also holds in the quantum domain and is experimentally confirmed in a plethora of phenomena from atomic spectra to particle acceleration. In the context of quantum information, it is the reason why the behavior of entanglement in relativity differs significantly from the standard, non-relativistic theory Gingrich and Adami (2002); Peres et al. (2002); Friis et al. (2010); Palge and Dunningham (2015); Barr et al. (2023).
In this paper we propose an experiment to verify Wigner’s rotation in the regime of what are typically regarded as non-relativistic velocities at $2\cdot 10^{3}$ m/s in an implementation with slow neutrons. This is significant because it would provide direct evidence of Wigner’s rotation in the quantum domain in a system that can be carefully controlled and, because the experiment does not require relativistic speeds, it can be carried out in a laboratory. From a geometric point of view the experiment can be regarded as explicitly verifying the curvature of the relativistic velocity space—a feature that sets apart the relativistic and non-relativistic theories of spacetime Palge and Pfeifer (2023).
The article is organized as follows. We begin by summarizing the theoretical derivation of the Wigner rotation in Sec. II followed by the discussion of the experimental setup with slow neutrons in Sec. III and the conclusion in Sec. IV.
II Wigner rotation
Wigner’s rotation originates in the fact that the subset of pure Lorentz boosts does not form a subgroup of the Lorentz group Halpern (1968); Costella et al. (2001); Rhodes and Semon (2004). The combination of two boosts $\Lambda(v_{1})$ and $\Lambda(v_{2})$ is in general a boost and a rotation,
$\Lambda(v_{2})\Lambda(v_{1})=W[\alpha(v_{1},v_{2})]\Lambda(v_{3}),$ (1)
where, for massive spin-$1/2$ particles that we will focus on, $W(\alpha)\in\mathrm{SO}(3)$ is the Wigner rotation with angle $\alpha$ about the axis $\hat{n}=\hat{v}_{2}\times\hat{v}_{1}$ orthogonal to the plane specified by $v_{1}$ and $v_{2}$. Wigner rotation is a truly relativistic phenomenon both in the conceptual and quantitative senses. The latter means that in a simple scenario consisting of two boosts the rotation angle becomes significant at large velocities. For instance, if $v_{1,2}=0.5c$ and the velocities are orthogonal, then the rotation is $8$ deg, while for non-relativistic speeds $v_{1,2}=2\cdot 10^{3}\,$ m/s it is barely measurable at $\,\sim\!\!10^{-5}$ deg. However, one can achieve a measurable effect at low, non-relativistic velocities by noting another property of the Wigner rotation, namely, that it is cumulative. This is because Wigner’s rotation is ultimately a holonomy—a geometric phase that accumulates when a particle follows a path in the momentum space, in analogy to how Berry’s phase accumulates when a particle follows a path in the parameter space associated with the Hamiltonian Wilczek and Shapere (1989); Lyre (2014); Palge and Pfeifer (2023).
In the following, we will take the geometric approach Palge and Pfeifer (2023), which is based on the fiber bundle theory Simon (1983); Cisowski et al. (2022), and give an outline of how the Wigner angle can be calculated based on the idea that it is a geometric phase. We will then use the result to design an experiment that allows to measure the Wigner rotation of a particle at non-relativistic speeds.
From the intuitive point of view, in the fiber bundle picture we can think of the state of the massive relativistic spin-$1/2$ particle as a field on curved momentum space. When a particle is Lorentz boosted it follows a geodesic in this manifold. This means the spin of the particle is parallel transported along the geodesic. Because the momentum space is curved, the state of the particle in general changes non-trivially when it is parallel transported along a given path. Hence, the final state after completing parallel transport along a given path is generally not identical to the initial state. For closed paths, the transformation that maps the initial state to the final state is called a holonomy. The Wigner rotation is thus nothing but the holonomy which arises due to the curvature of the relativistic momentum space. We will next describe how to compute the holonomy transformation for a massive relativistic spin-$1/2$ particle.
In the fiber bundle theory, the quantum relativistic massive spin-$1/2$ particle is described by the spinor bundle $(\pi:E\rightarrow\mathscr{V}^{+}_{m},\mathscr{H}_{\mathbf{p}})$ where, using the sign convention $\eta\sim(+,-,-,-)$, $\mathscr{V}^{+}_{m}$ is the forward mass hyperboloid $\mathscr{V}^{+}_{m,x}=\{P\in T^{*}_{x}\mathbb{E}^{1,3}\;|\;\eta(P,P)=\eta^{\mu\nu}p_{\mu}p_{\nu}=m^{2},p_{0}>0\}\subset T^{*}_{x}\mathbb{E}^{1,3}\,$ and the typical fiber $\mathscr{H}_{\mathbf{p}}=\mathbb{C}^{2}$, for a more detailed account see Palge and Pfeifer (2023). Because the Minkowski space is flat, we can identify the mass hyperboloids at different points of spacetime and speak of a single momentum hyperboloid $\mathscr{V}^{+}_{m}$ in which the particle moves. Using the spherical polar coordinates, the intrinsic geometry of the momentum hyperboloid is described by the metric
$g=-\left(\tfrac{m^{2}}{E^{2}}\mathrm{d}\rho\otimes\mathrm{d}\rho+\rho^{2}\mathrm{d}\theta\otimes\mathrm{d}\theta+\rho^{2}\sin^{2}\theta\mathrm{d}\phi\otimes\mathrm{d}\phi\right)\,.$ (2)
In order to describe parallel transport, we compute the coefficients $\omega^{A}{}_{B}$ of the spin connection using the relation
$\omega^{A}{}_{Bi}=e^{A}{}_{k}e_{B}{}^{j}{}\Gamma^{k}{}_{ij}+e^{A}{}_{k}\partial_{i}e_{B}{}^{k}{}\,,$ (3)
where $e^{A}{}_{i}$ are components of an orthonormal coframe $\Theta^{A}=e^{A}{}_{i}\mathop{}\!\mathrm{d}z^{i}$ of the momentum space metric $g$. We can display $\omega$ as the matrix
$\omega$ $=\begin{bmatrix}0&-\frac{E}{m}\mathop{}\!\mathrm{d}\theta&-\sin\theta\,\frac{E}{m}\mathop{}\!\mathrm{d}\phi\\ \frac{E}{m}\mathop{}\!\mathrm{d}\theta&0&-\cos\theta\mathop{}\!\mathrm{d}\phi\\ \sin\theta\,\frac{E}{m}\mathop{}\!\mathrm{d}\phi&\cos\theta\mathop{}\!\mathrm{d}\phi&0\end{bmatrix}.$ (7)
We can likewise express the curvature of the Levi-Civita connection as a collection of $2$-forms $\Omega^{A}{}_{B}$ over $\mathscr{V}^{+}_{m}$ written as the matrix
$\Omega=$ $\begin{bmatrix}0&-\frac{\sqrt{E^{2}-m^{2}}}{Em}\mathop{}\!\mathrm{d}\rho\wedge\mathop{}\!\mathrm{d}\theta&-\frac{\sqrt{E^{2}-m^{2}}}{Em}\sin\theta\mathop{}\!\mathrm{d}\rho\wedge\mathop{}\!\mathrm{d}\phi\\ \frac{\sqrt{E^{2}-m^{2}}}{Em}\mathop{}\!\mathrm{d}\rho\wedge\mathop{}\!\mathrm{d}\theta&0&-\frac{(E^{2}-m^{2})}{m^{2}}\sin\theta\mathop{}\!\mathrm{d}\theta\wedge\mathop{}\!\mathrm{d}\phi\\ \frac{\sqrt{E^{2}-m^{2}}}{Em}\sin\theta\mathop{}\!\mathrm{d}\rho\wedge\mathop{}\!\mathrm{d}\phi&\frac{(E^{2}-m^{2})}{m^{2}}\sin\theta\mathop{}\!\mathrm{d}\theta\wedge\mathop{}\!\mathrm{d}\phi&0\end{bmatrix}.$ (11)
The state space of the particle is given as the space of square integrable sections $\psi:\mathscr{V}^{+}_{m}\to\mathbb{C}^{2}$ of the spinor bundle with a suitable representation of the group $\mathrm{SU}(2)$. Formally, the spinor bundle is the associated bundle to the principal spin bundle whose structure group $\mathrm{Spin}(3)\cong\mathrm{SU}(2)$ is the double cover of $\mathrm{SO}(3)$.
In order to describe parallel transport of spinors, we need the spinor connection. This is induced by the Levi-Civita connection via lifting the $\mathrm{SO}(3)$ connection to the spin connection, and then generating the spinor connection. We get
$\omega_{s}=-\frac{i}{2}\left(\frac{E}{m}\mathop{}\!\mathrm{d}\theta\,\sigma^{3}-\frac{E}{m}\sin\theta\mathop{}\!\mathrm{d}\phi\,\sigma^{2}+\cos\theta\mathop{}\!\mathrm{d}\phi\,\sigma^{1}\right)\,.$ (12)
Using the spinor connection, we calculate the spinor curvature
$\Omega_{s}=\mathop{}\!\mathrm{d}\omega_{s}+\omega_{s}\wedge\omega_{s},$ (13)
which yields
$\Omega_{s}=$ $\frac{i}{2}\left(-\frac{\sqrt{E^{2}-m^{2}}}{Em}\mathop{}\!\mathrm{d}\rho\wedge\mathop{}\!\mathrm{d}\theta\,\sigma_{3}\right.$
$\phantom{AA}+\frac{\sqrt{E^{2}-m^{2}}}{Em}\sin\theta\mathop{}\!\mathrm{d}\rho\wedge\mathop{}\!\mathrm{d}\phi\,\sigma_{2}$
$\phantom{AA}-\left.\frac{E^{2}-m^{2}}{m^{2}}\sin\theta\mathop{}\!\mathrm{d}\theta\wedge\mathop{}\!\mathrm{d}\phi\,\sigma_{1}\right).$ (14)
Consider now a particle that follows circular path $C$ in the momentum space. This means the particle is moving with constant speed $V$, but constantly changing its direction. One can think of particle’s movement as undergoing infinitesimal parallel transports when it travels around the circular trajectory. Each small parallel transport corresponds to a small boost which gives rise to Wigner’s rotation, all of which accumulate when the particle has completed one orbital revolution. This corresponds to the famous case of Thomas precession, which is formally a holonomy associated with a circular path in the momentum space. The holonomy matrix can be calculated in terms of the spinor curvature $\Omega_{s}$
$\mathrm{Hol}\left(\omega_{s},C\right)=\exp\left(-\int_{D}\Omega_{s}\right),$ (15)
where $D$ is a disk with boundary $C$ and the matrix is a function of connection $\omega_{s}$ and path $C$. In our case, the path is a circle with radius $\rho_{0}\in\mathscr{V}^{+}_{m}$, where $\rho_{0}=mV/\sqrt{1-V^{2}}$. It is the boundary of the disc $D$ given by $\theta=\pi/2$, $\rho\in(0,\rho_{0})$ and $\phi\in(0,2\pi)$. Using (II) with $\mathop{}\!\mathrm{d}\theta=0$, the integral
$-\int_{D}\Omega_{s}$ $=-\int_{0}^{\rho_{0}}\int_{0}^{2\pi}\mathop{}\!\mathrm{d}\rho\mathop{}\!\mathrm{d}\phi\ \frac{i}{2}\frac{\rho}{\sqrt{m^{2}+\rho^{2}}m}\ \sigma_{2}$ (16)
$=-i\sigma_{2}\pi\left(\gamma(V)-1\right)\,.$ (17)
By substituting into (15) we obtain the matrix
$\mathrm{Hol}(\alpha)=\exp\left(-i\frac{\alpha}{2}\sigma_{2}\right)=\begin{pmatrix}\cos(\alpha/2)&-\sin(\alpha/2)\\ \sin(\alpha/2)&\cos(\alpha/2)\end{pmatrix}\,,$ (20)
where we use the argument $\alpha$ to emphasize that this is an $\mathrm{SU}(2)$ rotation matrix with an angle
$\alpha(V)=2\pi\left(\gamma(V)-1\right)$ (21)
which represents the holonomy transformation of the spin state after the particle has completed one revolution around $C$.
We will next describe an experiment that allows to measure Wigner’s rotation using slow neutrons.
III Experiment
[…]
$\mathop{\left|\nearrow\right\rangle}=\mathrm{Hol}(\alpha)^{N}\mathop{\left|\uparrow\right\rangle}=\mathrm{Hol}(N\alpha)\mathop{\left|\uparrow\right\rangle},$ (22)
thus producing the total rotation angle $N\alpha$ [^1] Note that the last equality in (22) is simply obtained by using the explicit form of the holonomy matrix (20) for the case under consideration.
A measurement $M$ carried out on the final spin state shows the difference between the directions of the initial and final spins. Note that since we assumed that the initial spin lies in the plane of the circle or, equivalently, in the plane of the boosts around the circle, the particle undergoes maximal Wigner rotation.
An interesting possibility for realizing the experiment are low-energy neutrons. They have been used extensively to probe fundamental concepts of quantum physics Hasegawa and Badurek (1999); Sponar et al. (2010b); Demirel et al. (2015); Sponar et al. (2021). Let us analyze the magnitude of the effect in a setup similar to the one in Fig. 1 which involves low energy neutrons traveling at $V=2\cdot 10^{3}$ m/s with the mean lifetime $\tau_{n}=880$ s or about $15$ minutes. Let us assume a ring-like guide with radius $r$ can be manufactured which makes neutrons follow a circular path. This can be achieved by using material whose Fermi (pseudo) potential $V_{F}$ is such that if neutrons approach the surface of the guide at grazing incidence with small angle $\theta$, they are reflected by the potential of the material. When dealing with low-energy neutrons, Fermi potential is the most important quantity that characterizes neutron interaction with the material Sponar et al. (2021). Reflection occurs if the kinetic energy $E$ corresponding to the normal component of the velocity is smaller than the Fermi potential of the material. In other words, for total reflection $E$ must satisfy the Fermi-Zinn condition $E\sin^{2}\theta<V_{F}$. For instance, for a typical material Ni the angle $\theta<0.1$ deg. It is important to note that, because the Fermi potential is spin independent, the reflection does not change the spin state Boothroyd (2020).
Assuming the circular neutron guide has radius $r$, the particle makes $V/2\pi r$ revolutions in unit time. Multiplying this with the duration $t$ of the experiment and the angle $\alpha(V)$ per one revolution given in (21), we obtain the total rotation
$\alpha_{T}(V,r,t)=\frac{V}{2\pi r}\cdot t\cdot\alpha(V).$ (23)
Plot of $\alpha_{T}(V,r,t)$ for slow neutrons with $t=10\tau_{n}$ as well as longer times is shown in Fig. 2.
Figure 2: Dependence of the total Wigner rotation $\alpha_{T}$ on the radius $r$ of circle for a particle with $V=2\cdot 10^{3}\,$m/s, and duration of experiment $t$ is $10\tau_{n}$ (blue), $20\tau_{n}$ (yellow), $30\tau_{n}$ (green) and $40\tau_{n}$ (red).
A total spin rotation of about $1.1$ deg can be observed when the duration of experiment is $150$ mins and the guiding ring radius is $2$ cm. This is close to the precision of $1$ deg achievable in a standard configuration. [^2] Better precision has been obtained in an experiment in Hasegawa et al. (2012). Since the total rotation is proportional to the duration of experiment, the effect will be twice as large at about $2.2$ deg for experiments that last twice as long, and even larger for longer times. Note that at $150$ mins the duration of the experiment exceeds neutron mean lifetime by ten times, which results in loss of signal. Although this is a restriction it is not a critical issue as long as the resulting signal can be detected. [^3] Free neutron decay is described by $N(t)=N(0)e^{-t/\tau_{n}}$, where $N(0)$ is the initial number of neutrons. This means after time period $t=2\tau_{n}$ about $13\%$ of the initial population remains, while only $0.0045\%$ is left after $10\tau_{n}$ has passed. Whether or not the final population of neutrons can be measured at the end of the experiment depends on the detector sensitivity and the initial size of the neutron population.
Since the total rotation angle $\alpha_{T}(V,r,t)$ is inversely proportional to radius $r$ of the guide and proportional to experiment time $t$, the total angle $1.1$ deg can be also achieved by reducing the radius of guide tenfold to $2$ mm while reducing experiment time tenfold to $15$ min. Because the latter is equal to the mean lifetime of neutrons the resulting signal would be significantly stronger. Along the same lines, if one would let the experiment run twice as long, $30$ mins, the resulting angle would double to $2.2$ deg. However, manufacturing a smaller guide might be considerably more difficult. In summary, we see that in principle the parameters can be controlled so different regimes of implementation are possible.
IV Conclusion
We have proposed an experiment that allows to verify Wigner’s rotation for massive spin-$1/2$ particles in the non-relativistic regime. Although we showed that spin rotation $1.1$ deg can be achieved in principle, the parameters determining the outcome are tunable, which allows for possible observation of a larger rotation angle.
We would like to emphasize that since slow neutrons are non-relativistic, the only relevant potential that characterizes the interaction between neutrons and the material is the Fermi potential, which does not change the direction of spin. The reason one can detect Wigner rotation at these low velocities rests on the cumulative character of the phenomenon: although each revolution around the circle at $V=2\cdot 10^{3}$ m/s produces a minuscule rotation of $\,\sim\!\!10^{-10}$ deg, this can be amplified to an observable $1.1$ deg simply by letting the particle complete a sufficient number of revolutions. The scheme given above is one possible way of implementing such an experiment.
V Acknowledgments
VP acknowledges partial support by the Estonian Research Council (Eesti Teadusagentuur, ETAG) through grant PSG489. CP is funded by the excellence cluster QuantumFrontiers of the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under Germany’s Excellence Strategy, EXC-2123 QuantumFrontiers, 390837967.